OpenStudy (anonymous):

for x=2^3 , Y=2^-3 and z=3^-1 , write each expression in simplest form with positive exponents a. xy

OpenStudy (anonymous):

\[xy=2^3\times 2^{-3}=2^{3-3}=2^0=1\]

OpenStudy (anonymous):

b. \[\frac{ x }{ y }\] c. \[x^{2}\] d. \[\sqrt[10]{z}\] e. \[\left( x + y \right)^2\] f.xyz

OpenStudy (anonymous):

\[\frac{x}{y}=\frac{2^3}{2^{-3}}=2^{3+3}=2^6\]

OpenStudy (anonymous):

how many of these you got?

OpenStudy (anonymous):

4 left :)

OpenStudy (anonymous):

\[x^2=(2^3)^2=2^6\]

OpenStudy (anonymous):

\[\sqrt[10]{z}=\sqrt[10]{3^{-1}}=3^{-\frac{1}{10}}=\frac{1}{\sqrt[10]{3}}\]

OpenStudy (anonymous):

\[(x+y)^2=x^2+2xy+y^2=2^6+2+2^{-6}\]

OpenStudy (anonymous):

and since \(xy=1\) we have \(xyz=z=3^{-1}\)